A coin is tossed 7 times. Let A be the event that the first toss is heads. Let B be the event that the fifth toss is heads.
Are A and B mutually exclusive? Why or why not?
Are A and B independent? Why or why not?

Answer by stanbon(75887)

A coin is tossed 7 times. Let A be the event that the first toss is heads. Let B be the event that the fifth toss is heads.
Are A and B mutually exclusive? Why or why not?
No because there are a slew of sequences of 7 tosses
that have heads on the 1st and heads on the 5th toss.
----------------------------------------------------
Are A and B independent? Why or why not?
Yes because the P(head) on 1st = 1/2 and P(head on 5th) = 1/2
and P(A nd B) = P(A)*P(B) = 1/4
===================================
cheers,
Stan H.